# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v2

@@ 5.4.3. Summary
In both studies, we observed a small decrease in cognitive load from the baseline to the treatment condition, but this difference was not significant according to the lab study results. Participants consistently perceived little difference in the tools with respect to temporal demand and frustration. However, their perceptions varied across the studies on whether Papers-to-Posts was more or less mentally demanding and whether it required more or less effort.
@@ 5.5.1. Lab Study - H4
We recorded two metrics for task completion time. One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post. The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post. As participants reached that point of comfort during most sessions, we focused on the former metric. Because this data was numeric within-subjects data for two conditions of one factor, we used a paired-samples t-test. Participants worked slightly faster with the baseline tool (M=44.32 minutes, SD=12.51) in comparison to Papers-to-Posts (M=45.99 minutes, SD=13.88) (Appendix 26(a)), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
(a) Lab Study
(b) Deployment Study
[…]
5.6. Fulfillment of Planning Design Goal (DG1)
We analyzed participants’ Likert-type responses to questions related to fulfillment of our first design goal, related to planning the blog post (Figure 8a-b,e-f).
5.6.1. Lab Study - Exploratory
Based on the median results, participants were more confident that all essential information was in the blog post (treatment: M=6.00, Q1=3.00, Q3=6.25; baseline: M=5.00, Q1=3.00, Q3=6.00) and found it easier to incorporate information from the paper (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.75, Q3=6.00) using Papers-to-Posts (Figure 8a-b). In addition, the interaction logs showed that participants utilized Papers-to-Posts’s DG1-inspired affordance to select relevant paper bullet points and paragraphs to incorporate in the blog post. All participants adjusted the LLM’s pre-selected bullet points to some degree. On average, participants added or removed 28 bullet points and 13 paragraphs. We note that some added bullet points come from automatic selection of bullet points used to generate a new section rather than manual selection. Looking at participants’ logs, when selecting points to generate text, only a few exclusively selected original-text paragraphs rather than bullet points. Participants also provided custom bullet points an average of 1.70 times.
5.6.2. Deployment Study - Exploratory
Looking at the median results, participants did not indicate an improvement in feeling confident that all essential information was covered with Papers-to-Posts (M=5.00, Q1=4.50, Q3=6.00) as opposed to the baseline (M=5.00, Q1=4.25, Q3=6.00). Given that the automatic method for parsing the paper PDF in the deployment study was likely to make more errors than the partially manual method for parsing papers in the lab study, it makes sense that participants would now see a reduced benefit from Papers-to-Posts’s paper outline in terms of ensuring all relevant information is included. Multiple participants under both conditions noted that the parsing of their papers was incomplete or flawed. However, participants still indicated an increased ease in incorporating information from the paper in the blog post with Papers-to-Posts (M=6.00, Q1=5.00, Q3=6.00) compared to the baseline (M=4.50, Q1=3.25, Q3=5.75). For instance, P14 commented, “I liked being able to check the boxes of the portions of the paper I wanted to include in each section.”
5.6.3. Summary
In both studies, participants indicated increased ease in incorporating paper content in the blog post with Papers-to-Posts. However, the two studies had different outcomes with respect to participants’ confidence in all essential information being covered in the blog post. While lab study participants indicated higher confidence with Papers-to-Posts  the deployment study participants did not. This may be because, in the deployment study, we did not include manual corrections to the paper parsing.
5.7. Fulfillment of Drafting and Revising Design Goals (DG2, DG3)
We analyzed participants’ Likert-type responses to questions related to fulfillment of our second and third design goals, related to drafting and revising the blog post (Figure 8c-d,g-h).
5.7.1. Lab Study - Exploratory
Participants’ Likert responses regarding DG2 and DG3 indicate that they found it easier to produce appropriate blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) and iterate on the blog post text (treatment: M=6.00, Q1=5.00, Q3=6.00; baseline: M=5.00, Q1=4.00, Q3=6.00) in the treatment condition (Figure 8c-d). Furthermore, the interaction logs for the treatment condition indicate that participants took advantage of DG2-based affordances for generating text and DG3-based affordances for modifying text (Appendix 25). For an average of 6.85 generations per treatment condition, participants often utilized custom instructions (5.70 times), selected paragraphs or bullet points (6.35 times), and included the context of prior blog post sections (6.00 times), though we note that these features were pre-filled or toggled on by default. In addition, they modified text 6.85 times on average, with the “condense” (1.85 times), “custom” (1.80 times), and “simpler terms” (1.40 times) modifications being the most common.
5.7.2. Deployment Study - Exploratory
There was no difference in the median for how well Papers-to-Posts (M=5.00, Q1=4.00, Q3=6.00) versus the baseline (M=5.00, Q1=5.00, Q3=6.00) facilitated production of appropriate text, unlike in the lab study (Figure 8e-h). Nevertheless, participants indicated that Papers-to-Posts (M=6.00, Q1=4.50, Q3=6.00) facilitated iterating on the blog post more so than the baseline (M=5.00, Q1=3.50, Q3=6.00).
5.7.3. Summary
In both studies, participants using Papers-to-Posts found it easier to iterate on the blog post. However, while lab study participants thought Papers-to-Posts improved their ability to produce appropriate text for the blog post, deployment study participants did not. Perhaps without time constraints, deployment study participants had more time to determine how to prompt the baseline tool or adjust its text in order to produce appropriate text.
5.8. Qualitative Analysis
@@
Across the two studies, participants’ survey and interview comments reflected the difference in editing power between the two tools.
5.8.1. Baseline
@@
Overall, participants using the baseline most commonly cited the initial draft (lab: 6/20, deployment: 8/18) as a helpful aspect of the tool. Thus, the most valued baseline feature was not helpful for editing power. With respect to the initial draft, multiple participants noted editing or adding to the initial draft rather than replacing it. For example, P20-dep explained, “It’s easier to critique than it is to generate,” and P26-dep shared, “It was very useful to have a straw-person of what the blog post could look like (generated by the LLM). I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.” A number of participants also noted the tool’s flexibility as a benefit (lab: 6/20, deployment: 4/18). For example, P18-lab shared, “I was able to provide instructions that fit my mental model of constructing a blog post and it being there as a ‘copilot’ is just what I need.” Flexibility in drafting text may help writers obtain the edits they want. However, the writer holds more responsibility in steering the LLM appropriately; P7-lab reflected, “As a CS student, I like this way more [than the treatment’s way of supporting writing], but I think it needs more thinking to use it creatively.
5.8.2. Papers-to-Posts’ Planning Features
@@
Participants often noted the selectable paper bullet points and paragraphs as a helpful feature of Papers-to-Posts (lab: 10/20, deployment: 14/19). Several described the outline as helpful for selecting relevant content to generate text. For instance, P15-lab noted, “I could rewrite entire sections or direct it with just a few bullets.” Seven participants mentioned that the outline was useful even just for reviewing the paper’s content or structure. P4-lab commented, “The outline was really helpful in providing a quick overview of what content I had written.” For two lab participants, the bullet points provided transparency around the origins of the initial LLM draft and generations. P12-lab explained, “I liked the bullet point features to see what was being fed in vs not.” By facilitating review and selection of content to include in the blog post, the paper outline may have contributed to participants’ increased editing power.
@@
That said, participants also specified difficulties related to the paper outline. For instance, two lab participants were overwhelmed with the number of bullet points. P9-lab described the outline as “really bloated and hard to navigate.” Meanwhile, three lab participants were confused or encumbered by aspects of utilizing the bullet points. For example, P3-lab “felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets.”
5.8.3. Papers-to-Posts’ Drafting and Revising Features
@@
Participants frequently mentioned the utility of Papers-to-Posts’s text modification abilities (lab: 12/20, deployment: 9/19). For example, P18-dep reflected, “The ‘MODIFY WRITING’ section is helpful. I like that the interface contains this particular section to support iterating on existing text.” All of the preset modification buttons (“simpler terms”, “condense”, “expand”, “more dramatic”, “less dramatic”) were specifically mentioned as helpful affordances at least once in the lab study alone. The most commonly noted one was the “condense” button. For example, P12-lab commented, “I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts. I found that helpful.” Three participants commented that they desired more preset modification buttons. For instance, P16-lab wanted a “more technical” button, while P14-dep wanted “more or less casual” options. P7-lab then added, “Or even let user to add their favorite rewrite button.” Interestingly, P2-lab found the preset modification buttons useful for understanding the tool’s capabilities: “It was nice to know what options the tool already knew of, and would perhaps have success with. For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’. I would have stuck to more basic things like ‘please summarize the following’.” Several participants (lab: 5/20, deployment: 10/19) also noted the usefulness of at least one aspect of the customizable instructions for either generating or modifying text using Papers-to-Posts. P9-shared, “…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” Thus, Papers-to-Posts’s affordances for text drafting and revision may have contributed to participants’ increased editing power.
@@
Participants also discussed difficulties that they encountered in drafting and revising writing with Papers-to-Posts. Four lab participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts, contrasting with the baseline’s flexibility that participants appreciated. As an example, two wanted assistance with checking for redundancies in the blog post. The other two wanted feedback on the blog post writing. Regarding the affordances for modifying text, 8/20 lab and 5/19 deployment participants noted a difficulty. Four lab participants raised confusion about how or when to use certain affordances. As an example, P6-lab was unsure of “when to use which one of the six [modification] options.” Another recurrent issue noted by five lab participants was a modification affordance not working as desired. For P20-lab, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned. For P12-lab, modifications were sometimes too exaggerated, while for P16-lab, a custom modification to make writing more technical was too subdued. A couple deployment participants also noted that modifying text sometimes introduced errors. Future work may address these points of restriction and confusion noted by the participants using Papers-to-Posts.
@@ 6. Discussion
Based on our results, we reflect on design implications for other mixed-initiative tools for research-paper summarization and long-form scientific communication.
@@ 6.1. Post-Editing is Not Enough
Both Papers-to-Posts and the baseline tool included initial LLM-generated drafts generated with powerful LLMs. In the lab study, they were GPT-4 and GPT-4-32k respectively; in the deployment study, both conditions used GPT-4-0125-preview. Nevertheless, while the participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so. This suggests that writers may benefit from additional support other than an initial draft when writing research-paper blog posts.
6.2. Increasing Editing Power with Plan-Draft-Revise Workflow
With additional features to support planning, drafting, and revising the blog post, Papers-to-Posts increased writers’ editing power (i.e., change in the blog post per minute or writing action). One reason participants may have demonstrated increased editing power is that we provided an interactive research-paper outline to help them plan content to include in their blog post. Indeed, when using Papers-to-Posts versus the baseline tool, the median participant found it easier to incorporate information from the paper in the blog post. In their qualitative responses, participants also consistently noted the interactive paper outline as one of the most useful features. We therefore note that presenting LLM-generated interactive outlines of research papers shows promise in making the planning step of writing their summaries easier. Another reason participants may have demonstrated increased editing power is that we provided pre-filled yet customizable LLM instructions for generating and modifying the blog post text. In alignment, the median participant found it easier to iterate on the blog post while using Papers-to-Posts. Several participants commented on the usefulness of different aspects of Papers-to-Posts’s preset yet customizable LLM instructions for drafting and revising the blog post text. Thus, we observe that preset LLM instructions scaffolded for customization may support users in drafting and revising long-form scientific communication.
Lowering the temporal cost and writing cost of making an edit, Papers-to-Posts facilitates writers’ iteration on their blog posts, which we saw may be helpful for creating a blog post that satisfies them in a time-constrained setting like the lab study. Iteration may also help the writers to reflect on the blog post text (Zhou and Sterman, 2024). On the other hand, participants often described the baseline as more flexible and allowing them more freedom. Thus, utilizing the two tools in combination may prove useful. Papers-to-Posts may be used when the writer desires suggestions for how to edit their blog post, and the baseline tool (or no LLM tool) could be used when the writer has a specific edit in mind or wants to give themselves creative space to contemplate their own edits (Buschek, 2024).
@@ 7. Limitations and Future Work
Most participants were PhD students in computer science, who may benefit from raising awareness of their early-career work through blog posts; future work may investigate how more senior researchers utilize a tool like Papers-to-Posts. Future studies can also examine how researchers in other fields, who are generally less familiar with LLMs and may have different traditional paper formats, may benefit from a tool like Papers-to-Posts. Furthermore, Papers-to-Posts was designed to support blog posts that summarize a paper, but there are a variety of scientific blog post formats that future work could support (e.g., narrative or opinion-based formats). Future work may additionally investigate how different aspects of the input paper such as length, recency, and author position affect the user experience and outcome.
Papers-to-Posts itself has several limitations. For one, multiple deployment study participants commented on how the automatic paper parsing makes mistakes and misses parts of the paper text. The tool also does not support incorporating figures and tables in the blog post. In addition, future research could work to make the connection between the outline and paper text clearer (e.g., highlight blog post text to reveal related paper text). Moreover, how detailed the paper outline should be remains uncertain. While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming. Lastly, future work may look into how writers’ expertise may be utilized in supporting their blog post writing; a tool for Q-and-A research-paper summarization, for example, utilizes the researcher’s background to personalize its output (Lim et al., 2024).
8. Ethical Considerations
It is well-known that LLMs may hallucinate information included in their generated text (Bubeck et al., 2023). We mitigate this risk by focusing on researchers summarizing their own work, which they should know well. Multiple participants identified inaccuracies or exaggerations in generated text, which they could then correct. However, a few participants raised concerns about catching errors, suggesting a need for better support in detecting them. Another ethical concern is whether using an LLM-powered tool to write paper blog posts could lead to the loss of important details. Reflecting on the baseline, P3-lab raised this point: “…it seems like the easier you make it, i.e., the better the initial summary, the less likely I am to myself remember those things that happen to be missing.” Nevertheless, the median participant found it easier to incorporate information from the paper with Papers-to-Posts, so this work provided one step in the direction of reducing loss of important details. Furthermore, while an LLM-powered tool for writing research-paper blog posts may contribute to making science more accessible, it may raise concerns about decreasing paper readership. However, blog posts neither provide the same amount of information nor go through peer review. Papers still hold their own value, and blog posts can act as catalysts for paper reading. Finally, researchers may not acknowledge the use of AI in writing their blog posts, which could mislead readers, particularly if the writer has not properly verified the content. Many participants who shared their blog posts chose to disclose its use, but to provide better transparency, future versions of Papers-to-Posts may highlight the provenance of the output writing (e.g., by visualizing the writer’s interaction with the tool (Naimul Hoque et al., 2023)).
@@ 9. Conclusion
We introduce Papers-to-Posts, an LLM-powered tool for writing research-paper blog posts that follows the Plan-Draft-Revise workflow. We validated that Papers-to-Posts supports these steps through two user studies comparing how researchers write about their own papers with Papers-to-Posts versus a strong baseline (LLM blog post draft and free-form LLM prompting). In a within-subjects lab study (N=20 participants), Papers-to-Posts significantly increased writers’ editing power (i.e., change in initial blog post within a given amount of time), significantly increased satisfaction with the output blog post, and did not increase cognitive load. In a more realistic setting, a deployment study (N=37 blog posts, 26 participants) also showed that Papers-to-Posts resulted in greater editing power in terms of both time and writing actions, without decreasing blog post satisfaction or increasing cognitive load. We hope that this work informs future research and tools related to research-paper summarization and long-form scientific communication.

### v3

@@ 5.6.3. Summary
In both studies, we observed little difference in cognitive load from the baseline to the treatment condition. Participants consistently perceived little difference in the tools with respect to mental demand, temporal demand, and frustration. However, their perceptions varied across the studies in terms of effort.
@@ 5.7.1. Lab Study - H4
We recorded two metrics for task completion time. One was the amount of time participants needed during the session before they would be comfortable publicly publishing the blog post. The other was the perceived amount of time participants would need after the session before they would be comfortable publicly publishing the blog post. As participants reached that point of comfort during most sessions, we focused on the former metric. The mean participant worked slightly faster with the baseline tool (M=44.32 minutes, SD=12.51) in comparison to Papers-to-Posts (M=45.99 minutes, SD=13.88) (Appendix 26(a)), but this difference was not significant (Paired-Samples t-Test, t(19)=0.55, p=n.s.).
5.8. Qualitative Analysis - DG1 and DG2
@@
Across the two studies, participants’ survey and interview comments reflected the difference between the two tools in terms of supporting understanding (DG1) and control (DG2) of what content is versus is not included in the blog post.
@@
Overall, participants using the baseline tool most commonly cited the initial draft (lab: 6/20, deployment: 8/18) as a helpful aspect of the tool. Thus, the most valued baseline feature was not useful for understanding or controlling what content was omitted from the detailed summary. With respect to the initial draft, multiple participants noted editing or adding to the initial draft rather than replacing it. For example, P20-dep explained, “It’s easier to critique than it is to generate,” and P26-dep shared, “It was very useful to have a straw-person of what the blog post could look like (generated by the LLM). I mostly used the draft as a place to edit and steer in the direction I wanted the post to take.” Without an affordance to help them consider all the possibilities of content to include in the blog post, the participants seem to have focused more on augmenting the content already selected by the tool.
@@
With respect to Papers-to-Posts, participants often noted the selectable paper bullet points and paragraphs as a helpful feature (lab: 10/20, deployment: 14/19). Several described the outline as helpful for selecting relevant content to generate text. For instance, P15-lab noted, “I could rewrite entire sections or direct it with just a few bullets.” Seven participants mentioned that the outline was useful even just for reviewing the paper’s content or structure. P4-lab commented, “The outline was really helpful in providing a quick overview of what content I had written [in the paper].” For two lab participants, the bullet points provided transparency around the origins of the initial LLM draft and generations. P12-lab explained, “I liked the bullet point features to see what was being fed in vs not.” Thus, participants’ interviews reflected how Papers-to-Posts’ outline affordance facilitated review and selection of content to include in the blog post.
@@
That said, participants also specified difficulties related to the outline mechanism. For instance, two lab participants were overwhelmed with the number of bullet points. P9-lab described the outline as “really bloated and hard to navigate.” Meanwhile, three lab participants were confused or encumbered by aspects of utilizing the bullet points. For example, P3-lab “felt stressed by choosing between the entire paragraph at hand and the (presumably) AI-generated summary bullets.” Future work may investigate how to make the outline mechanism more usable.
5.9. Qualitative Analysis - DG3
Across the two studies, participants’ survey and interview comments reflected the difference between the two tools in the tradeoff between supporting flexibility and providing scaffolding for alignment with academic blog post guidelines (DG3).
A number of participants noted the baseline tool’s flexibility as a benefit (lab: 6/20, deployment: 4/18). For example, P18-lab shared, “I was able to provide instructions that fit my mental model of constructing a blog post and it being there as a ‘copilot’ is just what I need.” Flexibility in drafting text may help writers obtain the edits they want. However, the writer holds more responsibility in steering the LLM appropriately; P7-lab reflected, “As a CS student, I like this way more [than the treatment’s way of supporting writing], but I think it needs more thinking to use it creatively.
@@ 5.9. Qualitative Analysis - DG3
Regarding Papers-to-Posts, participants frequently mentioned the utility of its text modification abilities, which were designed with the guidelines in mind (lab: 12/20, deployment: 9/19). For example, P18-dep reflected, “The ‘MODIFY WRITING’ section is helpful. I like that the interface contains this particular section to support iterating on existing text.” All of the preset modification buttons (“simpler terms”, “condense”, “expand”, “more dramatic”, “less dramatic”) were specifically mentioned as helpful affordances at least once in the lab study alone. The most commonly noted one was the “condense” button. For example, P12-lab commented, “I sometimes put the modify text through multiple rounds of expand and contract, focusing on different parts. I found that helpful.” Three participants commented that they desired more preset modification buttons. For instance, P16-lab wanted a “more technical” button, while P14-dep wanted “more or less casual” options. P7-lab then added, “Or even let user to add their favorite rewrite button.” Interestingly, P2-lab found the preset modification buttons useful for understanding the tool’s capabilities: “It was nice to know what options the tool already knew of, and would perhaps have success with. For example, I would probably not have gone to a blank-box tool and said, ’please make the following more dramatic’. I would have stuck to more basic things like ‘please summarize the following’.” Several participants (lab: 5/20, deployment: 10/19) also noted the usefulness of at least one aspect of the customizable instructions for either generating or modifying text using Papers-to-Posts. P9-shared, “…being able to include explicit instructions for the model to generate text from was helpful in being able to control the information in the text that it generated.” Thus, participants’ comments indicate that Papers-to-Posts’s preset yet flexible LLM instructions for generating academic blog post text provided utility.
@@
Participants also discussed difficulties that they encountered in devising instructions to Papers-to-Posts. Four lab participants expressed wanting to take specific actions that were not explicitly supported by Papers-to-Posts, contrasting with the baseline’s flexibility that participants appreciated. Two wanted assistance with checking for redundancies in the blog post, and the other two wanted feedback on the blog post writing. Regarding the affordances for modifying text, 8/20 lab and 5/19 deployment participants noted a difficulty. Four lab participants raised confusion about how or when to use certain affordances. As an example, P6-lab was unsure of “when to use which one of the six [modification] options.” Another recurring issue noted by five lab participants was a modification affordance not working as desired. For P20-lab, setting the desired length of the modified writing to one paragraph still resulted in three, as the back-end prompts are not guaranteed to work as planned. For P12-lab, modifications were sometimes too exaggerated, while for P16-lab, a custom modification to make writing more technical was too subdued. A couple deployment participants also noted that modifying text sometimes introduced errors. Future work may address these points of restriction and confusion noted by the participants using Papers-to-Posts.
@@ 6. Discussion
Based on our results, we reflect on design implications for other mixed-initiative tools for detailed long-document summarization.
@@ 6.1. Post-Editing is Not Enough
Both Papers-to-Posts and the baseline tool included initial drafts generated with powerful LLMs. Nevertheless, while participants had the option to simply post-edit the initial draft and refrain from interacting further with the LLM, only two baseline and two treatment participants across the lab and deployment studies did so. This suggests that writers may benefit from additional support other than an initial draft when writing detailed summaries of long-documents.
6.2. Increasing Understanding and Control of Content Selection for LLM-Generated Detailed Summaries
We presented a novel mechanism to support understanding and controlling the content of detailed long-document summaries– an interactive LLM-generated outline of the long document. This mechanism was implemented in Papers-to-Posts. Compared to a strong baseline, Papers-to-Posts made incorporating research-paper content in a blog post summary easier for researchers. Furthermore, under time constraints, researchers were more satisfied with their blog posts written using Papers-to-Posts, particularly with respect to their content coverage. Researchers also found it easier to iterate on their blog posts with Papers-to-Posts. Although the scaffolded and preset LLM instructions for adhering to academic blog post guidelines may have also facilitated iteration, researchers consistently noted the interactive paper outline as one of the most useful features of the tool. In alignment with increased ease of iteration, researchers experienced increased editing power (i.e., change in the blog post per minute or writing action) with Papers-to-Posts. We therefore conclude that LLM-generated interactive outlines of long documents show promise for making detailed summaries easier to write.
On the other hand, researchers often described the baseline as more flexible and allowing them more freedom. Thus, utilizing the two tools in combination may prove useful. A tool like Papers-to-Posts may be used when the writer is unsure of what content to include in their detailed long-document summary, and a tool like the baseline tool could be used when the writer already has a specific set of points in mind or wants to give themselves creative space to use the LLM in a variety of manners.
@@ 7. Limitations and Future Work
Most participants were PhD students in computer science, who may benefit from raising awareness of their early-career work through blog posts; future work may investigate how more senior researchers utilize a tool like Papers-to-Posts. Future studies can also examine how researchers in other fields, who are generally less familiar with LLMs and may have different traditional paper formats, may benefit from a tool like Papers-to-Posts. Future work may additionally investigate how different aspects of the input paper such as length, recency, and author position affect the user experience and outcome.
Papers-to-Posts itself has several limitations. For one, multiple deployment study participants commented on how the automatic paper parsing makes mistakes and misses parts of the paper text. The tool also does not support incorporating figures and tables in the blog post. Moreover, how detailed the paper outline should be remains uncertain. While several participants appreciated the bullet points, a couple mentioned that the outline was overwhelming. In addition, Papers-to-Posts’ outline only had one level of bullet points, but future work could explore the effects of a hierarchical outline. Lastly, future work may look into how writers’ expertise may be utilized in supporting their blog post writing; a recent tool for Q-and-A research-paper summarization, for example, utilizes the researcher’s background to personalize its output (Lim et al., 2024).
@@ 8. Conclusion
We introduce interactive reverse source outlines, a novel mechanism for summarizing long-documents in detailed articles. The mechanism consists of an LLM-generated reverse outline of the source document with pre-selected bullet points for an initial draft of the detailed summary, which the user manipulates to control the content covered in the summary. We implement this mechanism in Papers-to-Posts, an LLM-powered tool for writing research-paper blog posts. We validated that Papers-to-Posts helps users to understand and control their detailed summaries’ content through two user studies: a within-subjects lab study (N=20 participants) and a between-subjects deployment study (N=37 blog posts, 26 participants). The studies compared how researchers write about their own papers with Papers-to-Posts versus a strong baseline (LLM blog post draft and free-form LLM prompting). In the time-constrained within-subjects study, Papers-to-Posts significantly increased participants’ satisfaction with their final blog posts, especially in terms of the content covered. Furthermore, across both studies, participants found incorporating document content in their blog posts easier and demonstrated increased editing power. We hope that this work informs future research and tools related to human-LLM long-document summarization.